OpenStudy (boxman61):

need help simplifying... u=x^2 2u+3=0

OpenStudy (unklerhaukus):

\[2x^2+3=0\]

OpenStudy (boxman61):

so far i have... 2u=-3 u=-3/2 x^2 = -3/2 \[x=\pm \sqrt{-3/2}\]

OpenStudy (unklerhaukus):

\[2x^2=-3\] \[x^2=-3/2\] \[x=\sqrt{-1\times 3/2}\] \[x=\sqrt{\small{3/2}} \quad i\]

OpenStudy (boxman61):

thats what i thought but was not sure if it could simplify further...thank you

OpenStudy (unklerhaukus):

yeah put the ±in front of the last two lines i have used \[\sqrt{-1}=i\]